# Bowen and Wilson Image Class Translation Distance

### v1

@@ 4.1 Bone Marrow Smear Cytology
To accommodate more than two classes, we switched from a CycleGAN to a StarGAN.
@@
Utilizing the Bone Marrow Cytology in Hematologic Malignancies dataset, we achieved 95% training accuracy, and a 92% test accuracy in a 3-class identification task, predicting classes using the smallest translation distance. This showcases the method’s potential on larger tasks. This was a marginal improvement over the traditional CNN, which achieved an 89% test accuracy.
To get a better breakdown of class performance we evaluated the confusion matrices of both networks(Fig. 7 Although, the overall accuracy was close, there were differences in the class accuracies. Erythroblast class accuracy was similar (95% EfficientNet vs 96% StarGAN) with the StarGAN having a slight edge. The StarGAN also had a 3% higher accuracy for the lymphocytes and also was able to improve blast accuracy by 10% over the EfficientNet.
Blasts and lymphocytes are more likely to be mistaken for each other by both models. The EfficientNet had more blast false negatives and lymphocyte false positives than the StarGAN. Although, the StarGAN decreased the amount of false positives and negatives, the same underlying trend remains that blasts and lymphocytes are more likely to be mistaken for each other compared to erythroblasts.
@@
In image samples in Fig. 8, we can see some morphology changes that occurred for each of the images to fit into each class. The top row shows an image of a blast, that was labeled correctly by the image2image network, but misclassified by the CNN. The original blast is darkly stained, and it is difficult to distinguish the nucleus from the cytoplasm. The translation to a blast (in-class) retains these characteristics. The translation to an erythoblast resulted in a centered round nucleaus surrounded by a lighter-stained cytoplasm, while retaining the overall cell. Lastly, the translation to a lymphocyte changed the hue to a darker purple.
@@
The middle row of Fig. 8 shows an erythroblast that was correctly identified by both networks. Translation to an erythrocyte (in-class) retains the shape of cell and nucleus, and the contrast between cytoplasm and nucleus. Translation to a blast resulted in a loss distinction between the nucleus and cytoplasm, similar to the blast in the top row. Translation to a lymphocyte expanded the nucleus.
@@
The bottom row of final image panel shows a lymphocyte with a lighter-staining cytoplasm than the previous two examples. Translation to a lymphocyte (in-class) retains these characteristics. Translation to a blast image enlarged the nucleus and darkened the cytoplasm, consistent with the blast shown in the top row. Translation to an erythroblast, on the other hand shrank the nucleus and maintaned a lighter cytoplasm.
In these examples, the StarGAN followed by a translation distance classifier was able to correctly classify the images. The translation distances mirror showcases the classification decision, with the smallest distance being the classification. The CNN was able to classify erythroblast and lymphocyte correctly, but was unable to classify the blast.
We also highlight some image translation failure cases. Fig. 9 shows a correctly classified cell, however, there are hallucinations in the produced images, especially in the translation to blast. These hallucinations alter the image in such a way that the image no longer looks realistic. This is important in showing where the network might be overgeneralizing or is undertrained, but we note that the in-class translation still was achieved with lower translation distance than the other classes.
@@ 5.1 Bone Cytology
Expanding to six classes, the image2image network was able to achieve 90% testing accuracy (91.96% with a custom logistic regression classifier on translation distance) by classifying the image using the smallest translation distance. The traditional end-to-end CNN, by comparison achieved only 86% accuracy. (Fig. 1) Individual class accuracy ranged from 86%-97%.(Fig. 9)
The concern from the 2-class and 3-class problems was the potential that the network would not be able to generalize well or be expanded to more classes. However, this experiment showed that accuracy was still able to remain high with an increased number of classes and samples.
@@
The original three cytology classes were more morphologically similar: blasts, erythroblasts, and lymphocytes all have large, darkly stained nuclei. Lymphocytes differ by having a thin rim of lighter-stained cytoplasm. The three added classes added more variety to the morphology in the available images. For instance, the samples of the segmented neutrophil tended to have a characteristic clear cell body (chromophobic cytoplasm) and segmented nuclei. As a result, the generated images for that class needed to either add or modify features to resemble that characteristic. In Fig. 10, the segmented neutrophil has more morphological changes than the other images. As a result, the translation distance is greater and it does not closely resemble the input image.
Despite high overall accuracy, and high class accuracy, the network still made some classification errors. In Fig.11, the network incorrectly classified a blast as a promelocyte. This was an error that the traditional CNN made as well. The end-to-end CNN in this case has a misleading confidence of 99.8% that the image is a promyleocyte. This high-confidence prediction makes it hard to question the decision, but in this case, it is an overconfident and incorrect decision. The promyelocyte had the lowest translation distance and was selected as the class. However, a quick visual inspection shows that the hypothetical blast and promyelocyte both look visually similar to the input blast image, and both have small translation distances, indicating correctly, a much lower confidence than the end-to-end CNN suggests.
For the 6-class experiment, the translation from the identity mapping can be better visualized. This principle on identity loss in CycleGANs 10, emphasizes that an image should not need to adjusted if it stays within its own class. In Fig. 9, showcases images that were correctly classified. For the segmented neutrophil image, the identity mapping preserves the characteristics of the original images, whereas the other classes modify the image. Additionally, even where the image is incorrectly classified using translation distance alone in Fig. 9, images translated into their own domains preserve original characteristics. In this panel, other images do not modify the image either, which leads to closer overall translation distances between classes. Classes that did show alterations had much higher translations distances and were less likely to be considered as classifications initially.
@@
Finally, we note that conducting the same experiment with PyTorch’s Random Weighted Sampler to solve the class imbalance problem and achieves comparable accuracies to these experiments.
@@ 5.2 Translation Distance Classifiers
In the original experiments with a CycleGAN, we applied an SVM for the final classification. Adjusting the SVM, we were able to achieve high accuracy. However, in the multi-class classification, the translation distance was better able to reflect class differences. As a result, although an SVM could improve classification results, translation distance alone could be used for the final decision.
@@
| Lowest Translation Distance | 91.03 | 90.07 |
@@
To demonstrate and determine the best accuracy, we tested a variety of simple classifiers. The translation distance classifiers all had similar accuracies, and their confusion matrices were very similar with slight differences in individual class accuracies. Translation distance alone does not use another network; instead, the smallest distance is selected as the predicted class. Adding a classifier can improve accuracy with a slight cost of interpretability allowing the networks to maximize accuracy. A custom logistic regression model with a softmax activation function is the most effective for the 6-class problem, however, most of the classifiers shown in Tab. 1 can achieve acceptable accuracy results.
Although overall accuracy was similar between the classifiers, there were some differences between class accuracies, both within the networks and compared the other classifiers. In Fig. 9), blasts have the lowest class accuracy at 86% and segmented neutrophils have the highest with 97%. Selecting different auxilary classifiers, results in slightly different class accuracies and ranges.
In addition to outperforming the overall accuracy of the traditional network, the custom logistic regression classifer resulted in higher individual class accuracies. The traditional network achieved only a 65% blast accuracy (Fig. 9), but the custom logistic classifier returned 65% (Fig.9). Similiarily, the lymphocyte had 71% and 87% accuracies, respectively.
@@ 6 Conclusion
Using image-to-image networks for classification is built on the same idea behind the identity loss in CycleGANs 10–an image translated into its own domain/class should not require any changes. Conversely, images translated into domains/classes to which they do not belong require more changes. Thus by comparing visually the source and generated images, or quantifying image translation distances (e.g. ) it is possible to determine the class of a source image . This method reduces each input image into a point in translation distance space, , which can be analyzed for clusters and trends. Remarkably, in the two class experiment with melanoma, without having been explicitly trained to do so, the network described here grouped pigmented lesions in a similar manner to a dermatologist’s call on whether to perform a biopsy (see Fig. 5 and discussion). Furthermore, the overlap between benign lesions that had to be biopsied and melanomas raises the question of whether it is even possible to distinguish some types of benign lesions from melanomas. While the drive for improved accuracy often assumes all benign lesions are distinguishable from melanomas (and vice-versa) from a dermoscopy photograph alone, this is just not the case, as documented by earlier investigations in the limitations of dermoscopy 30. We note that other networks and training approaches can in principle be used, and may provide better results, e.g. contrastive unpaired translation 31, DRIT 32, cyclic autoencoders 33, transformers 34, and stable diffusion models 35.
Unlike existing image dimension reduction methods, e.g. principal component analysis 36, 37, convolutional autoencoders 38, stochastic neighbor embedding 39, etc., the dimensions of this new translation distance space are straightforward and interpretable as the amount of change an image must undergo to fit into a given class. The translation distances are independently significant and can be used as features in downstream classifiers, whose inputs and weights are then much easier to interpret than end-to-end classifiers.
The image2image framework was able to achieve comparable accuracies, even outperforming a traditional CNN. Furthermore, it provided greater insights into how the network was making decisions. In the bone cytology experiment, classification and visual errors could be more easily inspected. The baseline CNN classifier, an EfficientNetB3, is one that resulted from extensive architecture hyperparameter tuning, and also made use of transfer learning. The image translation networks, CycleGAN and StarGAN, had none of those advantages, relying on small identity loss adjustments. This suggests the possibility that with further development, image-to-image networks may achieve better accuracy on classification tasks.
The network struggled to perform without an adjustment for class imbalance though, as some groups were over represented. Selecting an equal number of samples from each class during training solves this problem. However, it does limit the number of images seen. To increase the samples during training, weighted sampling allows the traditional 80/20 split to be used. The class distribution problem can be solved either way.
The translation networks are trained on the task of creating realistic images that can deceive the discriminator. Using the translation distance to classify images, adds to the functionality of the already robust discriminator within the image2image network. The training and testing results of the classification aspects, are secondary metrics of an already trained network. Consequently, the network retains comparable training and testing accuracies even when trained on a reduced dataset was selected to mitigate the class-imbalance problem. CNNs have been observed struggle to apply their performance outside the training set. Therefore, the StarGAN’s ability to achieve comparable accuracy in a larger test set is promising.
With this approach, translation distances can be used for classification or can be analyzed with a simple classifier (e.g. an SVM, a linear classifier or any number of neural network approaches can be used). The resulting generated images can be visually inspected to drive more informed decisions and uncover insights into the training dataset. For example, consider the artifact of vignetting (darkening at the edge of images). A human observer would overlook this as irrelevant to detecting melanoma, but a training dataset that exhibits certain artifacts more frequently in one class or another would easily mislead a conventional classifier. An image-to-image approach has the potential to reveal the severity of artifacts on a case-by-case basis: an artifact’s presence in the source and both generated images indicates a robust translation result, whereas an artifact’s absence in one of the generated images indicates it was mistaken for a defining class-specific feature. Such mistakes may result from having artifacts that are unequally distributed between training set classes, as we uncovered with vignetting, skin tone, and scalebar markings in the melanoma versus benign dataset. Such insights from applying image-to-image networks to difficult classification tasks will hopefully inform efforts to produce better training datasets and in turn, improve outcomes with conventional classification CNNs.
@@
Finally, the image2image framework can help bridge the interpretability gap between technologies such as (AI, and microscopy techniques) with their users, clinicians. While the clinician-in-the-training-loop approach yields impressive improvements in accuracy and robustness against artifacts 15, our approach is less labor intensive and provides clinicians an opportunity to examine visual differences between classes at inference (point of care). Both approaches are complementary and add value to AI processing and evaluation of medical imaging by working alongside, rather than to replace clinicians, and have potential for both improving diagnoses and in training human observers to recognize subtle differences between classes. Furthermore, the use of image2image for classification can reveal biases, morphological similarities and differences and show new patterns in the data, creating new lines of inquiry.
@@ 7 Acknowledgments
We thank Priya Wolff (University of Colorado School of Medicine) and Yevgeniy Seminov (Mass. General Brigham) for helpful conversations and suggestions, and Bill Carpenter (Colorado State University) for technical support with computational resources. Experiments were carried out on the Riviera cluster (Data Science Research Institute, Colorado State University). This work was funded by the Boettcher Foundation through a Boettcher Collaboration Grant and a Boettcher Educational Enrichment Grant.

### v2

@@ 4.1 Bone Marrow Smear Cytology
Utilizing the subset of Bone Marrow Cytology in Hematologic Malignancies, we achieved 95% training accuracy, and a 92% test accuracy in a 3-class identification task, predicting classes using the smallest translation distance. This was a marginal improvement over the baseline CNN, which achieved an 89% test accuracy. Confusion matrices of both networks are shown in Fig. 12, showing more pronounced improvements, especially in the blast category, which saw a 10% accuracy improvement over the EfficientNet. Although blasts and lymphocytes continue to be difficult to resolve, I2I-CT translation distance cut the number of these misclassifications by roughly a factor of two compared to the baseline CNN.
@@
A selection of images correctly classified simply by smallest translation distance is shown in Fig. 13. Note that in each of these cases, the StarGAN primarily modified the cell at the center of the image, mostly leaving the others alone. The top row shows a blast that was misclassified by the baseline CNN but correctly identified by smallest translation distance. The original is darkly stained, and it is difficult to distinguish the nucleus from the cytoplasm. The in-class (blast) hypothetical retains these characteristics. The translation to an erythoblast resulted in a centered round nucleus surrounded by a lighter-stained cytoplasm, while retaining the overall cell shape. Lastly, the translation to a lymphocyte changed the overall hue to a darker purple with a thin, lighter rim.
@@
The middle row of Fig. 13 shows an erythroblast that was correctly classified by both approaches. Translation to an erythroblast (in-class) retains the shape of cell and nucleus, along with the contrast between cytoplasm and nucleus. Translation to a blast resulted in a loss distinction between the nucleus and cytoplasm, similar to the blast in the top row. Translation to a lymphocyte expanded the nucleus, leaving a thin, lighter rim.
@@
The bottom row of Fig. 13 shows a lymphocyte with a lighter-staining cytoplasm than the previous two examples. Translation to a lymphocyte (in-class) retains these characteristics. Translation to a blast image enlarged the nucleus and darkened the cytoplasm, consistent with the blast shown in the top row. Translation to an erythroblast, on the other hand shrank the nucleus and maintained a lighter cytoplasm.
A translation failure case is shown in Fig. 14. There are hallucinated artifacts in the produced images, especially in the translation to blast. These hallucinations alter the image in such a way that the image no longer looks realistic. However, as with the failure cases in the pigmented lesions dataset (Fig. 11), in-class translation retained the smallest distance and this error did not interfere with correct classification.
@@ 5.1 Bone Cytology
Expanding to six classes, the image2image network was able to achieve 90% testing accuracy by classifying the image using the smallest translation distance, and 91.96% with a custom logistic regression classifier on translation distance vectors. Individual class accuracy ranged from 86%-97% (Fig. 19). The traditional end-to-end CNN, by comparison achieved only 86% overall accuracy (Table 1).
@@
Figure 15 shows an example image (plasma cell) correctly identified by the smallest translation distance. In the 3-class case, each cell type was morphologically similar: blasts, erythroblasts, and lymphocytes all have large, darkly stained nuclei. Lymphocytes differ by having a thin rim of lighter-stained cytoplasm. In the 6-class case, the added cell types had more morphological variety. For instance, the samples of the segmented neutrophil tended to have a characteristic clear cell body (chromophobic cytoplasm) and segmented nuclei. As a result, the generated image for that class has more morphological changes than the other images.
Figure 16 shows an example image (blast) misclassified as a promelocyte by minimum translation distance. The baseline CNN misclassified this image as well, but with an alarmingly misleading confidence of 99.8%. The advantage of the I2I-CT is readily seen here, as a quick visual inspection shows that the hypothetical blast and promyelocyte both look visually similar to the input blast image, and both have small translation distances, indicating correctly, a much lower confidence (roughly, a coin toss between blast and promyelocyte) than the end-to-end CNN asserts.
Additional panels of images and I2I-CT hypotheticals are shown in Figs. 17 and 18. Figure 17 shows a selection of images correctly classified by smallest translation distance. As expected from the use of an identity loss, in-class translations (highlighted by yellow outlines) are visually most similar to the originals in terms of morphology, color, and texture. The most prominent morphological difference that can be seen is in the nuclei of segmented neutrophils versus other types; often, the network alters this morphology as needed, but can be seen to struggle, especially in translating the segmented neutrophil to an erythroblast and a plasma cell. By contrast, Figure 18 shows a selection of images misclassified by smallest translation distance. Typically, the in-class hypothetical (yellow outline) is still visually similar to the original, and translation distances are a close call, which indicates uncertainty. The exception is the segmented neutrophil (in-class , whereas ). Yet this image appears to be an anomaly in that there is no central cell in the field of view; as noted, the network appears to have learned for the most part to alter the central cell, leaving the others untouched.
@@
Finally, we note that conducting the same experiment with a random weighted sampler to solve the class imbalance problem and achieves comparable accuracies to these experiments.
@@ 5.2 Translation Distance Classifiers
To further improve translation distance classification, we tested a variety of simple classifiers and compared against classification by the smallest translation distance (Tab. 1). These all had similar accuracies, with a logistic regression model with a softmax activation function performing the best (91.96% test accuracy versus 86.07% for the baseline end-to-end CNN).
@@
| Smallest Translation Distance | 91.03 | 90.07 |
@@
Figure 19 shows the confusion matrix for the translation distance classifier using logistic regression for probabilities. Figure 20 shows the confusion matrix for the baseline end-to-end CNN. The highest for both was for segmented neutrophils at 97.7% for the translation distance classifier and 97.3% for the end-to-end CNN. The worst class accuracy for both was for blasts at 86% for the translation distance classifier and 65%, markedly worse for the end-to-end CNN.
6 Discussion
Using image2image networks for classification is built on the same idea behind the identity loss in CycleGANs [48]: an image translated into its own domain/class should not require any changes. Conversely, images translated into domains/classes to which they do not belong require more changes. Thus by comparing visually the source and generated images, or quantifying image translation distances it is possible to determine the class of a source image . This method reduces each input image into a point in translation distance space, , which can be analyzed for clusters and trends. In this proof-of-concept work, we found (mean average difference) to suffice, in spite of its well-known limitations compared to other methods that better quantify perceptual differences [3, 44]. This work used a CycleGAN for 2-class scenarios and a StarGAN for multi-class scenarios, but we note that other networks and training approaches can in principle be used, and may lead to better results, e.g. contrastive unpaired translation [27], DRIT [22], cyclic autoencoders [19], transformers [46], and stable diffusion models [30].
As with prior work in counterfactuals [11], especially on the melanoma images [11], we found visual inspection revealed the CNN to be inadvertently sensitive to features that a domain expert would ignore (e.g. scalebars, background pigmentation, vignetting). The predominance of such features, or of generated hallucination artifacts may be helpful in flagging potentially misclassified images– see Figs. 9, 10, and 11. As with prior work in ANT-GAN on MRI images [38], we were able to train a CycleGAN to transform between benign and malignant pigmented lesions. Then, instead of subtracting the source image from the generated image to highlight pixels responsible for the differences, we analyzed translation distance vectors.
Translation distance, unlike existing image dimension reduction methods (e.g. principal component analysis [45, 24], convolutional autoencoders [6], stochastic neighbor embedding [41], which was used to compare distributions of real versus hypothetical MRI brain images in ANT [38], etc.) are straightforward and interpretable. Each dimension represents the amount of change an image must undergo to fit into a given class. Ideally, the true class will show the least alteration. As we showed, the translation distance vectors have significance both for discovering trends and clusters in the data and as features in downstream classifiers.
On the pigmented lesions task, translation distances revealed a natural clustering along the lines of whether or not a dermatologist called for a biopsy, rather than between malignant and benign (Fig. 7). This may be an interesting alternative training target for triaging pigmented lesions. Furthermore, the overlap between benign lesions that had to be biopsied and melanomas raises the question of whether it is even possible to distinguish some types of benign lesions from melanomas. While the drive for improved accuracy often assumes all benign lesions are distinguishable from melanomas (and vice-versa) from a dermoscopy photograph alone, this is just not the case, as documented by earlier investigations in the limitations of dermoscopy [36]. Notably, even though pigmented lesion classifiers are known for poor generalization to external data sources [43, 11], the region occupied by melanomas in translation distance space (trained from ISIC 2020 data) holds up under different datasets (ISIC 2016 and 2017). Further, we went beyond prior works by showing an extension to more than two classes, and demonstrated use of the translation distance vector itself as a feature vector that can be directly used for classification.
Compared to a conventional end-to-end CNN classifier, we found that classification based on simple -dimensional translation distance vectors achieved comparable accuracies. Two-class cases resulted in slightly worse accuracy, but 3- and 6-class cases resulted in consistently higher accuracy compared to a conventional end-to-end CNN. This could be due to differences between CycleGAN () and StarGAN (), differences between the datasets, or higher-dimensionality features of cases. Regardless, this is remarkable, considering, as with nearly all such end-to-end approaches, the EfficientNet’s classifier operates on a feature space from the convolutional front end having dimensions. By comparison, -dimensional translation distance vectors are a highly efficient representation apropos to classification. Further, our results from multiple years of ISIC dermoscopy data suggest translation distance features to be robust to out-of-distribution images. In addition, our hyperparameter and architecture turning was limited to small adjustments to and . Considering the baseline EfficentNet had the advantages of extensive architecture hyperparameter tuning [39] and transfer learning, its failure to significantly and consistently outperform translation distance classifiers is remarkable. This suggests the possibility that with further development, I2I-CT may enable better accuracy than currently possible. Another promising avenue is to extend the ANT-GAN approach of feeding class-specific hypotheticals along with the original image as additional input channels to a conventional classifier, which showed significant improvements in the context of MRI and brain tumor classification [38]. Finally, I2I-CT provided greater insights into how the network was making decisions compared to a black box classifier. In the bone cytology experiment, classification and visual errors could be more easily inspected.
It is also worth mentioning limitations of the I2I-CT approach. First, its complexity – although I2I-CT produces highly compact feature vectors, it requires multiple CNNs (generator(s) and discriminator(s)) – compared to a simple end-to-end classifier CNN, may limit applicability in cases of large images (e.g. gigapixel whole-slide images) or 3D datasets. Second, the use of CNNs for the generator(s) and discriminator(s) mean the same fundamental limitations of these architectures for end-to-end classifiers apply here. Depending on the application, though, the ability to inspect the class-specific hypotheticals and translation distance distributions may yield insights not readily apparent with end-to-end classifiers, and translation distance classifiers might in some cases yield better accuracy.
@@ 7 Conclusion
In conclusion, image2image networks applied to translation between classes (I2I-CT) opens up unique opportunities for understanding datasets, inspecting classifier decisions, and potentially improving upon end-to-end classifiers. Generated class-specific hypothetical images can reveal training dataset artifacts that can bias or mislead conventional classifiers. The impact of such artifacts can be observed on a case-by-case basis: an artifact’s presence in the source and both generated images indicates a robust translation result, whereas an artifact’s absence in one of the generated images indicates it was mistaken for a defining class-specific feature. When multiple class-specific hypotheticals look similar to the input image, this can indicate uncertainty even in cases where an end-to-end CNN misclassifies the images with high confidence. Further, each image can be reduced to a simple length- vector of translation distances, a low-dimensional representation that can be used for visualizing image distributions, identifying natural clusters in the data, and as an input to simple classifiers. Compared to the -dimensional feature vectors found in modern end-to-end CNN classifiers, translation distance vectors are a highly efficient representation apropos to classification. Our results on bone marrow cytology images even show the possibility of better accuracy with translation distance classifiers than end-to-end CNNs, though it remains to be seen whether this holds more broadly in other datasets.
@@
Finally, the image2image framework can help bridge the interpretability gap between technologies such as (AI, and microscopy techniques) with their users, clinicians. While the clinician-in-the-training-loop approach yields impressive improvements in accuracy and robustness against artifacts [11], our approach is less labor intensive and provides clinicians an opportunity to examine visual differences between classes at inference (point of care). Both approaches are complementary and add value to AI processing and evaluation of medical imaging by working alongside, rather than to replace clinicians, and have potential for both improving diagnoses and in training human observers to recognize subtle differences between classes. Finally, the use of image2image for classification may yield insights into difficult classification tasks and datasets that will hopefully inform efforts to produce better training datasets and in turn, improve outcomes with conventional classification CNNs.
8 Disclosures
The authors declare that there are no financial interests, commercial affiliations, or other potential conflicts of interest that could have influenced the objectivity of this research or the writing of this paper.
9 Code, and Data Availability
The code is available at the GitHub repository: https://github.com/mikylab/image2image, a fork of the original StarGAN repository[9].
The Apples to Oranges data [48] is publicly available at https://efrosgans.eecs.berkeley.edu/cyclegan/datasets/.
The ISIC Skin Lesion dataset [31] is available at https://challenge2020.isic-archive.com.
The Bone Marrow Cytology Dataset [25] is publicly available at https://www.cancerimagingarchive.net/collection/bone-marrow-cytomorphology_mll_helmholtz_fraunhofer/.
@@ 10 Acknowledgments
We thank Priya Wolff (University of Colorado School of Medicine) and Yevgeniy Seminov (Mass. General Brigham) for helpful conversations and suggestions, Veronica Rotemberg (Memorial Sloan Kettering Cancer Center) for discussions on lesion difficulty, and Bill Carpenter (Colorado State University) for technical support with computational resources. Experiments were carried out on the Riviera cluster (Data Science Research Institute, Colorado State University). This work was funded by the Boettcher Foundation through a Boettcher Collaboration Grant and a Boettcher Educational Enrichment Grant.